# The Role of the Schwinger Effect in Superradiant Axion Lasers

### v2

@@ I Introduction
In models that include an axion, rotating black holes are therefore expected to produce a dense axion cloud via superradiance. It has been proposed [23, 24, 25, 26, 27] that, in certain regions of parameter space, these axions can undergo stimulated decay into photons, creating a powerful laser, referred to in [23] as a BLAST (Black hole Lasers powered by Axion SuperradianT instabilities). It has been observed [23, 25] that these BLASTs are powerful enough that the Schwinger effect may become significant. The Schwinger effect refers to a quantum mechanical, non-perturbative process whereby a strong electromagnetic field produces electron-positron pairs. Plasma effects are significant in the dynamic of BLASTs [24, 25, 27], primarily because the plasma imparts photons with an effective mass, which eventually makes axion decay energetically impossible, thereby shutting off the laser. [23] speculated that plasma effects would result in the axion laser shutting off and then restarting in a periodic fashion, although they did not explore the underlying dynamics in detail. [25, 26] performed numerical simulations that seemed to validate the prediction of periodic bursts, whereas the simulations in [27] found that laser-like emission occurs not in bursts, but smoothly; however, all of these simulations assumed an electron-ion plasma acquired via accretion, as opposed to an electron-positron plasma acquired via the Schwinger effect. This paper therefore aims to explore the role of Schwinger production, specifically, in axion lasers.
@@ II BLASTs without Schwinger production
where $K\sim\mathcal{O}(1)$ is a model-dependent factor. Due to the quasi-bound nature of the $2p$-state, axions will not leave the $2p$-cloud except when disturbed by outside forces. However, a photon produced inside the $2p$-cloud will naturally exit the $2p$-cloud at the speed of light, producing an escape rate of
@@
[…]
@@
[…]
@@ III.1 Rate of Schwinger production
The Schwinger effect [39, 40] is a nonperturbative effect caused by vacuum decay in the presence of an electromagnetic field. The Schwinger effect can only occur if $E^{2}-B^{2}$ and $\vec{E}\cdot\vec{B}$ are nonzero; consequently, a single beam of light cannot induce the Schwinger effect. However, when multiple beams of light intersect, the interference of the two electromagnetic fields may satisfy the requirements for Schwinger production [41, 42]; this is an area of significant research interest, as it is hoped that the Schwinger effect may be observable at the focus of two lasers (for a review, see [43]). In the same way, Schwinger production may occur inside the $2p$-cloud, where a large number of photons are traveling in different directions.
For simplicity, we make two approximations. The first is that we ignore the magnetic field in calculating the rate of Schwinger production. This is common for research on the Schwinger effect at the intersection of laser beams, e.g. [41, 42, 43, 44, 45], for the reason that the Schwinger effect cannot occur in a pure magnetic field (i.e., zero electric field), whereas it can occur in a pure electric field (i.e., zero magnetic field). However, it should be noted that the contribution of the magnetic field can be significant in situations like the one we are analyzing, and it tends to reduce the rate of Schwinger production by 1-2 orders of magnitude [46]. The second approximation we make is that we neglect the effect of axions on the Schwinger effect; it has been shown that the presence of axions can enhance the Schwinger effect by several orders of magnitude [47]. Thus, our two approximations have opposite effects (increasing and decreasing the rate of Schwinger production), so that, while it is unclear whether our calculated pair production rate is an overestimate or an underestimate, it is likely accurate to the true value to within 1-2 orders of magnitude. The following should therefore only be understood as a rough order-of-magnitude estimate, which may be refined in future research.
In the case of a constant, uniform electric field, electron-positron pairs are created at a rate per unit volume of
$\frac{\textrm{d}n_{e^{+}e^{-}}}{\textrm{d}t}=\frac{e^{2}E^{2}}{4\pi^{3}}\sum_{n=1}^{\infty}\frac{1}{n^{2}}e^{-\pi\frac{E_{c}}{E}n},$ (19)
[…]
@@
In our case, the random decay of axions into photons results in a superposition of an enormous number of electric fields, so that the total electric field within the $2p$-cloud will be highly inhomogeneous and impossible to predict. However, the length- and timescales of these inhomogeneities will be the wavelength of the light, which is $\frac{2}{m_{\phi}}$. For all axion masses that are of interest as a dark matter candidate, this is far greater than the Compton wavelength. Therefore we may divide the $2p$-cloud into a number of patches, each with volume equal to the Compton volume $\lambda_{C}^{3}$, and in each patch we may treat the electric field as uniform and constant. (It should be noted that most research on the Schwinger effect at the intersection of laser beams is done under the assumption of high-frequency light, typically X-rays, and in such cases this approximation does not hold. We are taking advantage of the fact that the light in a BLAST is at a much lower frequency than what would be practical in an earth-based experiment.) Therefore, the total rate of Schwinger pair production throughout the $2p$-cloud is
$\frac{\textrm{d}N_{e^{+}e^{-}}}{\textrm{d}t}=\sum_{\textrm{patches}}\lambda_{C}^{3}\frac{e^{2}E^{2}}{4\pi^{3}}\sum_{n=1}^{\infty}\frac{1}{n^{2}}e^{-\pi\frac{E_{c}}{E}n}.$ (20)
@@
If one wished, one could now find the average value of the electric field in the $2p$-cloud, and substitute this value into the above equation. However, this approach would not produce an accurate estimate of the rate of Schwinger production. This is because of the Schwinger effect’s highly nonlinear nature, meaning that even slight increases to the electric field may result in a dramatic increase to the rate of pair production. This is significant in our case, because we have an electric field that is, effectively, randomly varying, and as such there will be some patches where the electric field strength is greater or less than the $2p$-cloud’s average field strength. These “hotspots” and “coldspots” contribute more than average and less than average, respectively, to pair production. If the Schwinger effect were linear, the hotspots and coldspots would cancel out when averaging over the patches, and therefore they could be ignored. However, the Schwinger effect’s nonlinear nature means that the surplus of pair production in the hotspots will be far greater than the deficit in pair production in the coldspots. Consequently, if we were to use the average value of $E$, we would significantly underestimate the rate of pair production. What we do instead is find the probability distribution of the electric field throughout the $2p$-cloud, $f(E)$. Our total rate of pair production will then be
$\frac{\textrm{d}N_{e^{+}e^{-}}}{\textrm{d}t}=V\int_{0}^{\infty}\textrm{d}Ef(E)\frac{e^{2}E^{2}}{4\pi^{3}}\sum_{n=1}^{\infty}\frac{1}{n^{2}}e^{-\pi\frac{E_{c}}{E}n}.$ (21)
@@
To find $f(E)$, we assume that the electric field components are normally distributed uncorrelated random variables with equal standard deviation, $\sigma_{\textrm{EM}}$. The electric field magnitude $E=\sqrt{E_{x}^{2}+E_{y}^{2}+E_{z}^{2}}$ then follows a chi distribution, with probability density
@@
$\sigma_{\textrm{EM}}$ may be found from the energy density, which in terms of the electromagnetic field components is
@@
Thus the rate of pair production is
$\frac{\textrm{d}N_{e^{+}e^{-}}}{\textrm{d}t}=\frac{3\sqrt{3}\pi e^{2}E_{c}^{5}V^{\frac{5}{2}}}{64\sqrt{m_{\phi}^{3}N_{\gamma}^{3}}}\sum_{n=1}^{\infty}n^{3}G^{30}_{03}\left(\frac{3\pi^{2}E_{c}^{2}V}{4m_{\phi}N_{\gamma}}n^{2}\begin{tabular}[]{|c}$-$\\ $-\frac{5}{2},-2,0$\end{tabular}\right)\equiv\Gamma_{\textrm{Schw}},$ (25)
where $G$ is the Meijer G-function. For every electron-positron pair produced, the electromagnetic field loses energy $2m_{e}$, resulting in the loss of photons at a rate of $\frac{4m_{e}}{m_{\phi}}\Gamma_{\textrm{Schw}}$.
We can see that Schwinger production is parameterized by $\frac{N_{\gamma}}{N_{\gamma}^{\textrm{Schw}}}$, where
@@
For $N_{\gamma}\lesssim N_{\gamma}^{\textrm{Schw}}$, $\Gamma_{\textrm{Schw}}$ is well approximated by
$\Gamma_{\textrm{Schw}}\approx\frac{e^{2}E_{c}^{2}V}{4\sqrt{3}\pi}\sum_{n=1}^{\infty}e^{-3\left(\frac{N_{\gamma}}{n^{2}N_{\gamma}^{\textrm{Schw}}}\right)^{\frac{-1}{3}}}\left(\frac{N_{\gamma}}{n^{2}N_{\gamma}^{\textrm{Schw}}}\right)^{\frac{1}{3}}.$ (27)
@@
We will find that, for our purposes, $N_{\gamma}$ is always less than $N_{\gamma}^{\textrm{Schw}}$, and therefore we may freely use this approximation.

### v3

@@ I Introduction
In models that include an axion, rotating black holes are therefore expected to produce a dense axion cloud via superradiance. It has been proposed [23, 24, 25, 26, 27] that, in certain regions of parameter space, these axions can undergo stimulated decay into photons, creating a powerful laser, referred to in [23] as a BLAST (Black hole Lasers powered by Axion SuperradianT instabilities). It has been observed [23, 25] that these BLASTs are powerful enough that the Schwinger effect may become significant. The Schwinger effect refers to a quantum mechanical, nonperturbative process whereby a strong electromagnetic field produces electron-positron pairs. Plasma effects are significant in the dynamic of BLASTs [24, 25, 27], primarily because the plasma imparts photons with an effective mass, which eventually makes axion decay energetically impossible, thereby shutting off the laser. [23] speculated that plasma effects would result in the axion laser shutting off and then restarting in a periodic fashion, although they did not explore the underlying dynamics in detail. [25, 26] performed numerical simulations that seemed to validate the prediction of periodic bursts, whereas the simulations in [27] found that laser-like emission occurs not in bursts, but smoothly; however, all of these simulations assumed an electron-ion plasma acquired via accretion, as opposed to an electron-positron plasma acquired via the Schwinger effect. This paper therefore aims to explore the role of Schwinger production, specifically, in axion lasers.
@@ II BLASTs without Schwinger production
where $K$ is a model-dependent factor. For the more general axion-like particle, $K$ is a free parameter, but for the QCD axion, which we examine in this paper, $K\sim\mathcal{O}(1)$. Due to the quasi-bound nature of the $2p$-state, axions will not leave the $2p$-cloud except when disturbed by outside forces. However, a photon produced inside the $2p$-cloud will naturally exit the $2p$-cloud at the speed of light, producing an escape rate of
@@
Superradiance requires $\alpha\lesssim.5$, and therefore we must have $m_{\phi}\gtrsim 10^{-8}\textrm{ eV}$ and $M_{BH}\lesssim 10^{-2}M_{\odot}$. This constraint restricts BLASTs to occur only around primordial Kerr black holes. Primordial black holes are expected to form with some small initial spin [28, 29, 30, 31, 32, 33], and they may spin up via a number of processes, including mergers [34], accretion [35, 36], hyperbolic encounters [37], and Hawking radiation [38]. Thus it is possible that the conditions for a BLAST might form in nature.
@@
$L^{\textrm{max}}=\frac{m_{\phi}}{2}N_{\gamma}^{\textrm{max}}\Gamma_{\gamma}=1.0*10^{43}\tilde{a}K^{-2}\left(\frac{\alpha}{0.037}\right)^{7}\left(\frac{m_{\phi}}{10^{-5}\textrm{ eV}}\right)^{-2}\left(\frac{\xi}{106}\right)\textrm{ erg/s}.$ (18)
@@ III.1 Rate of Schwinger production
The Schwinger effect [39, 40] is a nonperturbative effect caused by vacuum decay in the presence of an electromagnetic field. The Schwinger effect cannot occur when $E^{2}-B^{2}=\vec{E}\cdot\vec{B}=0$, and therefore a single beam of light cannot induce the Schwinger effect. However, when multiple beams of light intersect, the interference of the two electromagnetic fields may satisfy the requirements for Schwinger production [41, 42]; this is an area of significant research interest, as it is hoped that the Schwinger effect may be observable at the focus of two lasers (for a review, see [43]). In the same way, Schwinger production may occur inside the $2p$-cloud, where a large number of photons are traveling in different directions.
The Schwinger effect is driven not only by the electromagnetic field, but by the axion field as well [44, 45]. This is a consequence of the coupling between axions and electrons, which takes the form
$\mathcal{L}\supset\frac{g_{ae}}{2m_{e}}\bar{\psi}_{e}\gamma^{\mu}\gamma^{5}\psi_{e}\partial_{\mu}\phi,$ (19)
where $g_{ae}$ is the axion-electron coupling constant, $m_{e}$ is the electron mass, $\phi$ is the axion field, and $\psi_{e}$ is the electron bispinor. The presence of axions has the effect of enhancing the rate of pair production, compared to the Schwinger effect without axions; this is referred to as the axion-assisted Schwinger effect [44, 45]. This is an important consideration, however, for reasons that will be given momentarily, we do not include it in this paper, instead reserving it for future research. Therefore, the rate of Schwinger production given herein must be understood as a lower bound, as we are neglecting an effect that is known to enhance pair production. Put another way, the rate of Schwinger production that is used in this paper may be understood as the correct value only when $g_{ae}=0$; for $g_{ae}>0$, the rate of Schwinger production will be greater than what is used in this paper. Future research will be needed to determine the rate of Schwinger production when $g_{ae}>0$.
In the following subsection, we discuss the rate of Schwinger production assuming no axion assistance; after that we will discuss the difficulties presented by axion assistance. Finally, we present an alternative way in which the rate of pair production may be partially derived, in agreement with our result from the first subsection.
III.1.1 The Schwinger effect in Compton-volume patches
The Schwinger effect is typically calculated in the frame of reference in which the electric and magnetic field are parallel, which is guaranteed to exist except in the case where $E^{2}-B^{2}=\vec{E}\cdot\vec{B}=0$. The strength of the electric and magnetic fields in this frame of reference can be found by taking advantage of the Lorentz invariance of $E^{2}-B^{2}$ and $\vec{E}\cdot\vec{B}$:
$E_{\parallel}$ $=\sqrt{\frac{1}{2}\left(\sqrt{\left(E^{2}-B^{2}\right)^{2}+4(\vec{E}\cdot\vec{B})^{2}}+\left(E^{2}-B^{2}\right)\right)}$ (20)
$B_{\parallel}$ $=\sqrt{\frac{1}{2}\left(\sqrt{\left(E^{2}-B^{2}\right)^{2}+4(\vec{E}\cdot\vec{B})^{2}}-\left(E^{2}-B^{2}\right)\right)}.$ (21)
For a constant, homogenous electromagnetic field, the rate of Schwinger production is
$\frac{\textrm{d}n_{e^{+}e^{-}}}{\textrm{d}t}=\frac{e^{2}E_{\parallel}B_{\parallel}}{4\pi^{2}}\sum_{n=1}^{\infty}\frac{1}{n}\coth\frac{n\pi B_{\parallel}}{E_{\parallel}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}},$ (22)
where $E_{c}=\frac{m_{e}^{2}}{e}=1.3*10^{18}\textrm{ V/m}$ is the critical Schwinger field. Note that this increases with $E_{\parallel}$ but decreases with $B_{\parallel}$. In general, the Schwinger effect is highly dependent on the space- and time-dependence of the electromagnetic field, i.e. a nonconstant and/or nonuniform field will produce electrons and positrons at a very different rate than the figure given above. However, the effects of nonconstancy and nonuniformity are most pronounced when the electromagnetic field varies on length- and time-scales less than the Compton wavelength of an electron, $\lambda_{C}$ [46].
@@ III.1.1 The Schwinger effect in Compton-volume patches
In our case, the random decay of axions into photons results in a superposition of an enormous number of electromagnetic fields, so that the total electromagnetic field within the $2p$-cloud will be highly inhomogeneous and impossible to predict. However, the length- and time-scales of these inhomogeneities will be the wavelength of the light, which is $\frac{2}{m_{\phi}}$. For all axion masses that are of interest as a dark matter candidate, this is far greater than the Compton wavelength. Therefore we may divide the $2p$-cloud into a number of patches, each with volume equal to the Compton volume $\lambda_{C}^{3}$, and in each patch we may treat the electromagnetic field as uniform and constant. (It should be noted that most research on the Schwinger effect at the intersection of laser beams is done under the assumption of high-frequency light, typically X-rays, and in such cases this approximation does not hold. We are taking advantage of the fact that the light in a BLAST is at a much lower frequency than what would be practical in an earth-based experiment.) Therefore, the total rate of Schwinger pair production throughout the $2p$-cloud is
$\frac{\textrm{d}N_{e^{+}e^{-}}}{\textrm{d}t}=\sum_{\textrm{patches}}\lambda_{C}^{3}\frac{e^{2}E_{\parallel}B_{\parallel}}{4\pi^{2}}\sum_{n=1}^{\infty}\frac{1}{n}\coth\frac{n\pi B_{\parallel}}{E_{\parallel}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}.$ (23)
@@
If one wished, one could now find the average value of $E_{\parallel}$ and $B_{\parallel}$ in the $2p$-cloud, and substitute these values into the above equation. However, this approach would not produce an accurate estimate of the rate of Schwinger production. This is because of the Schwinger effect’s highly nonlinear nature, meaning that even slight increases to $E_{\parallel}$ or decreases to $B_{\parallel}$ may result in a dramatic increase to the rate of pair production. This is significant in our case, because we have an electromagnetic field that is, effectively, randomly varying, and as such there will be some patches where the electromagnetic field produces electron-positron pairs at a greater or lesser rate than would be predicted by the cloud-wide averages of $E_{\parallel}$ and $B_{\parallel}$. These “hotspots” and “coldspots” contribute more than average and less than average, respectively, to pair production. If the Schwinger effect were linear, the hotspots and coldspots would cancel out when averaging over the patches, and therefore they could be ignored. However, the Schwinger effect’s nonlinear nature means that the surplus of pair production in the hotspots will be far greater than the deficit in pair production in the coldspots. Consequently, if we were to use the average values of $E_{\parallel}$ and $B_{\parallel}$, we would significantly underestimate the rate of pair production. What we do instead is find the probability distribution of the electric and magnetic field strength throughout the $2p$-cloud, $f(E)$ and $f(B)$, and average over these. Our total rate of pair production will then be
$\frac{\textrm{d}N_{e^{+}e^{-}}}{\textrm{d}t}=V\int_{0}^{\pi}\frac{\textrm{d}\theta\sin\theta}{2}\int_{0}^{\infty}\int_{0}^{\infty}\textrm{d}E\textrm{d}Bf(E)f(B)\frac{e^{2}E_{\parallel}B_{\parallel}}{4\pi^{2}}\sum_{n=1}^{\infty}\frac{1}{n}\coth\frac{n\pi B_{\parallel}}{E_{\parallel}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}\equiv\Gamma_{\textrm{Schw}},$ (24)
where $\theta$ is the angle between the electric and magnetic fields (assumed to be uniformly distributed between 0 and $\pi$), and $E_{\parallel}$ and $B_{\parallel}$ are functions of $E$, $B$, and $\theta$. Note that $E$ and $B$ are the electric and magnetic field strengths in the axion cloud’s zero-momentum frame, whereas $E_{\parallel}$ and $B_{\parallel}$ are the electric and magnetic field strengths in a different frame of reference (that being the one in which the electric and magnetic fields are parallel).
@@
To find $f(E)$ and $f(B)$, we assume that the electromagnetic field components are normally distributed uncorrelated random variables with equal standard deviation, $\sigma_{\textrm{EM}}$. The electric field magnitude $E=\sqrt{E_{x}^{2}+E_{y}^{2}+E_{z}^{2}}$ then follows a chi distribution, with probability density
@@
An analogous expression for $f(B)$ may be found in the same way. $\sigma_{\textrm{EM}}$ may be found from the energy density, which in terms of the electromagnetic field components is
@@
We perform the integral in eq. 24 in Appendix A and show that
$\Gamma_{\textrm{Schw}}\approx\frac{e^{2}E_{c}^{2}V}{8\sqrt{3}\pi}\sum_{n=1}^{\infty}e^{-3\left(\frac{N_{\gamma}}{n^{2}N_{\gamma}^{\textrm{Schw}}}\right)^{\frac{-1}{3}}}\left(\frac{N_{\gamma}}{n^{2}N_{\gamma}^{\textrm{Schw}}}\right)^{\frac{1}{3}}$ (28)
where
@@
This approximation holds for $N_{\gamma}\lesssim.1N_{\gamma}^{\textrm{Schw}}$. We will find that, for our purposes, this is always the case, so that we may freely use this approximation for $\Gamma_{\textrm{Schw}}$.
III.1.2 The challenges of axion-assistance
In principle, the approach we’ve taken can easily be extended to account for axion-assistance. The axion field will vary in three ways with respect to time and space: it has a time-dependent phase factor $e^{-im_{\phi}t}$, its magnitude increases with time as a consequence of superradiant growth, and it is spatially inhomogeneous, which results from solving the Klein-Gordon equation in Kerr spacetime. These spatial inhomogeneities have lengthscale $\frac{2}{\alpha m_{\phi}}$ [13], and superradiant growth occurs over timescale $\frac{1}{2\Gamma_{s}}$. The shortest timescale is that which corresponds to the phase factor, $\frac{1}{m_{\phi}}$, and therefore this is the dominant form of axion assistance (since the axion couples to the electron via its derivative, $\partial_{\mu}\phi$). All of these length- and timescales are far larger than $\lambda_{C}$, and so, when we divide the axion cloud into Compton volumes, we may consider only the dominant term and approximate
$\phi=e^{-im_{\phi}t}\lvert\phi\rvert$. (30)
In principle, all we need do is derive an analog to eq. 22 for the case of $\phi=e^{-im_{\phi}t}\lvert\phi\rvert$, i.e. find the rate of Schwinger production when the electric field, magnetic field, and magnitude of the axion field are all held constant. We may then repeat the same procedure as before, inserting this analog of eq. 22 into eq. 24.
The challenge here lies in finding this analog to eq. 22. To the author’s knowledge, doing so analytically is not a solved problem in the literature; previous works on the axion-assisted Schwinger effect [44, 45] have been primarily numerical in nature. The prescription in these papers is to write a differential equation governing the time evolution of the Bogoliubov coefficients for the fermion, and then to numerically solve this differential equation. This yields the rate of pair production at a single point in momentum-space. To arrive at the total rate of pair production within a single Compton-volume patch (i.e., our analog to eq. 22), one would need to integrate over momentum space. Then, to find the total rate of pair production throughout the entire $2p$-cloud, one would need to integrate over $E$, $B$, and $\theta$, analogous to eq. 24. This would yield the rate of pair production in the $2p$-cloud, but only for a single value of $N_{\gamma}$ and $N_{\phi}$. Because $N_{\gamma}$ and $N_{\phi}$ will change over time, one would need to repeat this procedure as the axion laser evolves with time. The upshot is that numerically modeling the impact of the axion-assisted Schwinger effect on a superradiant axion laser would require an octuple integral (one integral to find the Bogoliubov coefficients, three over momentum space, three over the electric and magnetic fields, and one over time). As this is unlikely to be computationally feasible, it is necessary to determine an analytical expression to serve as our eq. 22-analog.
Deriving the rate of Schwinger production in the presence of both a constant electromagnetic field and an axion field that is constant up to a phase factor, is a sufficiently significant problem to deserve being a paper unto itself. For this paper, we defer that problem, neglecting axion assistance but noting that our expression for $\Gamma_{\textrm{Schw}}$ nevertheless has value as a lower bound. A more complete understanding of the Schwinger effect in BLASTs, applicable to the case of $g_{ae}>0$, will only be possible when the aforementioned eq. 22-analog has been derived.
III.1.3 An alternate perspective on pair production
In the preceding subsections, we approached this problem from the perspective of the classical electromagnetic field; we argued that, for sufficiently large $N_{\gamma}$, the field is sufficient to induce pair production at a significant rate. There is an alternative way of considering the problem, which is to focus on the individual photons present in the $2p$-cloud, produced by axion decay. This is an attractive lens through which to view the situation, as we know to a good approximation the spectrum and distribution of these photons (i.e., monochromatic, and uniformly distributed both in space and propagation direction). At the level of the individual photons, the Schwinger effect is caused by interactions between photons and a single electron-positron loop. It is an interesting question to see if we can derive our earlier results from the perspective of these specific particle interactions.
[47] calculates the rate at which $N$ photons, $K$ of which have positive helicity, will undergo a scattering interaction involving a single loop. They find that, within an electromagnetic field defined by the tensor $F_{\mu\nu}$, such a scattering interaction will take place at the rate
$\Gamma_{N,K}=-\frac{m_{e}^{4}}{8\pi^{2}}\left(\frac{2e}{m_{e}^{2}}\right)^{N}c\left(\frac{K}{2},\frac{N-K}{2}\right)\chi_{+}^{\frac{K}{2}}\chi_{-}^{\frac{N-K}{2}},$ (31)
where $c$ is a function defined in [47], and $\chi_{\pm}$ are related to the electromagnetic field:
$\chi_{\pm}=\frac{1}{8}F_{\mu\nu}F^{\mu\nu}\pm\frac{i}{8}F_{\mu\nu}\tilde{F}^{\mu\nu}.$ (32)
Without knowing the specific form of $F_{\mu\nu}$ (which will vary as a function of time and position), by considering the energy density of the electromagnetic field we may say that, for a cloud of photons with angular frequency $\omega$ and number density $n_{\gamma}$, we will have
$F_{\mu\nu}=\sqrt{n_{\gamma}\omega}f_{\mu\nu}(\vec{x},t),$ (33)
where $f_{\mu\nu}$ is a dimensionless function of time and space. The rate of scattering will therefore be
$\Gamma_{N,K}\propto m_{e}^{4}\left(\frac{e^{2}\omega}{m_{e}^{4}}n_{\gamma}\right)^{\frac{N}{2}},$ (34)
where the constant of proportionality will depend on $N$ and $K$, as well as time and position within the photon cloud. If $\frac{e^{2}\omega}{m_{e}^{4}}n_{\gamma}\ll 1$, $\Gamma_{N,K}$ will vanish for large $N$, and the rate of one-loop interaction will be negligible. Conversely, there exists some critical number density of photons above which pair production is significant, given by
$n_{\gamma}^{\textrm{crit}}\sim\frac{m_{e}^{4}}{e^{2}\omega}.$ (35)
Bearing in mind that $E_{c}=\frac{m_{e}^{2}}{e}$ and $\omega=\frac{m_{\phi}}{2}$, this reproduces, up to a constant of proportionality, eq. 29.